p=v.i but i=v/r so p=v^2/r so p is invesly proposnal to r

Grade:12th pass

2 Answers

yathartha gupta
6 years ago
P=vi but we know thati=V/R so we can do p=V^2RIn this way we see that p is inversely proportional TO RNICE ANSWER YOU KNOW
yathartha gupta
6 years ago
P=vi but we know thati=V/R so we can do p=V^2RIn this way we see that p is inversely proportional TO R . HENCE PROVED